A lead-acid battery uses a redox reaction in which lead(0) and lead(IV) are both converted to lead(II). This reaction is facilitated by the presence of sulfuric acid, H2SO4, as shown by the reaction

Pb+PbO2+2H2SO4→2PbSO4+2H2O

Part A Suppose that a fully charged lead-acid battery contains 1.50 L of 5.00 M H2SO4. What will be the concentration of H2SO4 in the battery after 2.70 A of current is drawn from the battery for 5.00 hours ?

Express your answer with the appropriate units.
Answer #1

H+ + 1e- ------> H

1 mol of H requires 1 mol of electron

1 mol of electron = 96485 C

So,1 mol of H requires 96485 C

let us calculate the charge passed:

t = 5.0 hr = 5.0*3600 s = 18000 s

time, t = 18000s

Q = I*t

= 2.7A * 18000s

= 48600 C

mol of H+ removed = 48600/96485 = 0.5037 mol

so,

mol of H2SO4 reacted = 0.5037 / 2 = 0.25185 mol

initially mol of H2SO4 = M*V = 5.00 M * 1.5 L = 7.5 mol

so,

mol of H2SO4 remaining = 7.5 mol -0.25185 mol = 7.248 M

So,

[H2SO4] = mol of H2SO4 / volume

= 7.248 mol / 1.50 L

= 4.83 M

Answer: 4.83 M